OpenStudy (anonymous):

Let f(x) = x2 - 3x - 7. Find f(-3). 11 8 -7 -10

OpenStudy (anonymous):

f(x) = x^2 -3x - 7 f(3) = 3^2 - 3(3) - 7 = ... f(3) simply means everywhere you see x, replace it with 3 t's f(-3), not f(3) f(-3) = -3^2 -3(-3) - 7 = 9 + 9 - 7 = 11 \ is c